What is the mass of hydrogen in 50.0 g of propyne, C3H4?
1
Expert's answer
2018-02-07T13:30:08-0500
.m(C3H4)=50 g
.m(H)=?
Mr(C3H4) = 3Ar(C) +4Ar(H)=3*12 +4=40
Amount of substance of propyne n(C3H4)= m/Mr=50/40=1.25 mol
.n(H)=4 n(C3H4)= 4*1.25= 5 mol
.m(H)= n(H)* Ar(H)= 5*1= 5g